OpenStudy (anonymous):

please help me substitution method 5x-y=12 9x-y=10

OpenStudy (anonymous):

alright so here's what you do. Get one of the equations in terms of another. For simplicity's sake we'll solve the first one for y

OpenStudy (anonymous):

subtract 5x from both sides so you get -y=-5x+12

OpenStudy (anonymous):

divide by -1 on both

OpenStudy (anonymous):

y=5x-12

OpenStudy (anonymous):

Plug (5x-12) into the second equation for y \[9x-(5x-12)=10\]

OpenStudy (anonymous):

9x-5x+12=10

OpenStudy (anonymous):

4x+12=10 4x+12-12=10-12 4x=-2 x=-0.5 or 1/2

OpenStudy (anonymous):

replace x in either equation with -0.5 or -1/2

OpenStudy (anonymous):

solve for y and you got it.